# A mathematical theory of topological invariants of quantum spin systems

### v2

@@ 3.1 Observables and derivations
The requirement that $X\cap Y\neq\varnothing$ is clear, since ${\mathsf{F}}^{X}$ and ${\mathsf{G}}^{Y}$ would commute otherwise. Suppose $Z\nsubseteq X\vee Y$. Then $Z^{\prime}:=(X\vee Y)\cap Z$ is a brick that is strictly contained in $Z$, so $[{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]^{Z}\in{{\mathfrak{d}}_{l}}^{Z}\cap{{\mathfrak{d}}_{l}}(Z^{\prime})=\{0\}.$
We make the following definitions for the next lemma. For any $U\subset{\mathbb{R}}^{n}$ and any brick $X$ write $\tilde{d}(X,U):=1+\operatorname{diam}(X)+d(X,U)$ and $\tilde{d}(X):=1+\operatorname{diam}(X)$.
Lemma 3.5.
For any $U\subset{\mathbb{R}}^{n}$, $Z\in{\mathbb{B}}_{n}$, and $k\geq 0$ we have
$\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}\tilde{d}(X,U)^{-k-4n-4}\tilde{d}(Y)^{-k-4n-4}\leq\frac{\pi^{8}16^{n}(n+1)^{4}3^{k}}{1296}\tilde{d}(Z,U)^{-k}.$
@@
Let $X,Y\in{\mathbb{B}}_{n}$ with $X\cap Y\neq\varnothing$ and let $Z\subset X\vee Y$ be a brick. Pick an arbitrary point $w\in X\cap Y$. We have
@@
$\leq 2\operatorname{diam}(X)+\operatorname{diam}(Y)+d(X,U),$
and so, since by Lemma 3.3 $\operatorname{diam}(Z)\leq\operatorname{diam}(X\vee Y)\leq\operatorname{diam}(X)+\operatorname{diam}(Y)$, we have
@@
It follows that for any $k>0$ we have
$\tilde{d}(Z,U)^{k}\tilde{d}(X,U)^{-k}\tilde{d}(Y)^{-k}$ $\leq 3^{k}$
@@
$\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}\tilde{d}(X,U)^{-k-4n-4}\tilde{d}(Y)^{-k-4n-4}\leq$
$\hskip 28.45274pt\leq 3^{k}\tilde{d}(Z,U)^{-k}\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}\tilde{d}(X,U)^{-4n-4}\tilde{d}(Y)^{-4n-4}$ (15)
It remains to bound the sum on the right-hand side. Fix an arbitrary $z\in Z$ and let $X,Y\in{\mathbb{B}}_{n}$ with $X\cap Y\neq\varnothing$ and $Z\subset X\vee Y$. Then by the second statement in Lemma 3.3 and the inequality $1+a+b\leq(1+a)(1+b)$ for $a,b\geq 0$ we have
@@
$\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}\tilde{d}(X,U)^{-4n-4}\tilde{d}(Y)^{-4n-4}\leq\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}(1+\operatorname{diam}(X))^{-4n-4}(1+\operatorname{diam}(Y))^{-4n-4}\\ \leq\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\end{subarray}}(1+\operatorname{diam}(X)+d(X,z))^{-2n-2}(1+\operatorname{diam}(Y)+d(Y,z))^{-2n-2}\\ \leq\left(\sum_{\begin{subarray}{c}X\in{\mathbb{B}}_{n}\end{subarray}}(1+\operatorname{diam}(X)+d(X,z))^{-2n-2}\right)^{2}\leq\frac{\pi^{8}16^{n}(n+1)^{4}}{1296}.$ (17)
Now we are ready to prove Proposition 3.6.
Proof of Proposition 3.6.
Notice that by Lemma 3.2 we have $\|{\mathsf{G}}\|_{V,k+2n+2}\leq\|{\mathsf{G}}\|_{{\mathbb{R}}^{n},k+2n+2}$ so without loss of generality we set $V={\mathbb{R}}^{n}$.
Let $Z\in{\mathbb{B}}_{n}$. By Lemmas 3.4 and 3.5 we have
$\sum_{X,Y\in{\mathbb{B}}_{n}}\|[{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]^{Z}\|$ $=\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}\|[{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]^{Z}\|$
@@
$\leq 2\|{\mathsf{F}}\|_{U,k+4n+4}\|{\mathsf{G}}\|_{{\mathbb{R}}^{n},k+4n+4}\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subset X\vee Y\end{subarray}}\tilde{d}(X,U)^{-k-4n-4}\tilde{d}(Y)^{-k-4n-4}$
$\leq C3^{k}\|{\mathsf{F}}\|_{U,k+4n+4}\|{\mathsf{G}}\|_{{\mathbb{R}}^{n},k+4n+4}\tilde{d}(Z,U)^{-k},$
with $C=\frac{\pi^{8}16^{n}(n+1)^{4}}{648}$. This proves that (11) is absolutely convergent and establishes the bound (12).
Next, let us prove the Jacobi identity. Let ${\mathsf{F}},{\mathsf{G}},{\mathsf{H}}\in{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$. We have
$[{\mathsf{F}},[{\mathsf{G}},{\mathsf{H}}]]^{W}$ $=\sum_{X,Y\in{\mathbb{B}}_{n}}\sum_{X^{\prime},Y^{\prime}\in{\mathbb{B}}_{n}}\left[{\mathsf{F}}^{X},\left[{\mathsf{G}}^{X^{\prime}},{\mathsf{H}}^{Y^{\prime}}\right]^{Y}\right]^{W}$ (18)
To show that this sum is absolutely convergent, we bound
$\sum_{X,Y\in{\mathbb{B}}_{n}}\sum_{X^{\prime},Y^{\prime}\in{\mathbb{B}}_{n}}\left\|\left[{\mathsf{F}}^{X},\left[{\mathsf{G}}^{X^{\prime}},{\mathsf{H}}^{Y^{\prime}}\right]^{Y}\right]^{W}\right\|$
@@
$\leq$ $4\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ W\subset X\vee Y\end{subarray}}\|{\mathsf{F}}^{X}\|\sum_{\begin{subarray}{c}X^{\prime},Y^{\prime}\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Y\subset X^{\prime}\vee Y^{\prime}\end{subarray}}\|{\mathsf{G}}^{X^{\prime}}\|\|{\mathsf{H}}^{Y^{\prime}}\|$
$\leq$ $C4\|{\mathsf{G}}\|_{{\mathbb{R}}^{n},8n+8}\|{\mathsf{H}}\|_{{\mathbb{R}}^{n},8n+8}\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ W\subset X\vee Y\end{subarray}}\|{\mathsf{F}}^{X}\|\tilde{d}(Y)^{-4n-4}$
$\leq$ $C4\|{\mathsf{F}}\|_{{\mathbb{R}}^{n},4n+4}\|{\mathsf{G}}\|_{{\mathbb{R}}^{n},8n+8}\|{\mathsf{H}}\|_{{\mathbb{R}}^{n},8n+8}\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ W\subset X\vee Y\end{subarray}}\tilde{d}(X)^{-4n-4}\tilde{d}(Y)^{-4n-4}$
$<$ $\infty$.∞ .
Here we used Lemma 3.5 in the second and fourth lines and $C$ is a constant depending only on $n$. Thus the sum (18) is absolutely convergent. In particular, using the fact that $\sum_{Y\in{\mathbb{B}}_{n}}[{\mathsf{G}}^{X^{\prime}},{\mathsf{H}}^{Y^{\prime}}]^{Y}=[{\mathsf{G}}^{X^{\prime}},{\mathsf{H}}^{Y^{\prime}}]$ we have the following absolutely convergent expression
$[{\mathsf{F}},[{\mathsf{G}},{\mathsf{H}}]]^{W}=\sum_{X,Y,Z\in{\mathbb{B}}_{n}}[{\mathsf{F}}^{X},[{\mathsf{G}}^{Y},{\mathsf{H}}^{Z}]]^{W}.$
It is then easy to check that the Jacobi identity for the sum follows from the Jacobi identity for each term. ∎
From Propositions 3.4 and 3.6 we immediately get
Corollary 3.1.
Suppose $U,V\in{\mathbb{R}}^{n}$.

### v3

@@ 2.1 Observables and derivations
The requirement that $X\cap Y\neq\varnothing$ is clear, since ${\mathcal{A}}^{X}$ and ${\mathcal{B}}^{Y}$ would commute otherwise. Suppose $Z\nsubseteq X\vee Y$. Then $Z^{\prime}:=(X\vee Y)\cap Z$ is a brick that is strictly contained in $Z$, so the ${{\mathfrak{d}}_{l}}^{Z}$-component of ${\mathcal{A}}^{X},{\mathcal{B}}^{Y}]$ is in ${{\mathfrak{d}}_{l}}^{Z}\cap{{\mathfrak{d}}_{l}}(Z^{\prime})=\{0\}.$
We make the following definitions for the next lemma. For any $U\subseteq{\mathbb{R}}^{n}$ and any brick $X$ write $\tilde{d}(X,U):=1+\operatorname{diam}(X)+d(X,U)$ and $\tilde{d}(X):=1+\operatorname{diam}(X)$.
Lemma 33.
For any $U\subseteq{\mathbb{R}}^{n}$, $Z\in{\mathbb{B}}_{n}$, and $k\geq 0$ we have
$\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subseteq X\vee Y\end{subarray}}\tilde{d}(X,U)^{-k-4n-4}\tilde{d}(Y)^{-k-4n-4}\leq\frac{\pi^{8}16^{n}(n+1)^{4}3^{k}}{1296}\tilde{d}(Z,U)^{-k}.$ (46)
@@
Let $X,Y\in{\mathbb{B}}_{n}$ with $X\cap Y\neq\varnothing$ and let $Z\subseteq X\vee Y$ be a brick. Pick an arbitrary point $w\in X\cap Y$. We have
@@
$\leq 2\operatorname{diam}(X)+\operatorname{diam}(Y)+d(X,U),$ (47)
and so, since by Lemma 31 $\operatorname{diam}(Z)\leq\operatorname{diam}(X\vee Y)\leq\operatorname{diam}(X)+\operatorname{diam}(Y)$, we have
@@
It follows that for any $k\geq 0$ we have
$\tilde{d}(Z,U)^{k}\tilde{d}(X,U)^{-k}\tilde{d}(Y)^{-k}$ $\leq 3^{k}$ (49)
@@
$\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subseteq X\vee Y\end{subarray}}\tilde{d}(X,U)^{-k-4n-4}\tilde{d}(Y)^{-k-4n-4}\leq$
$\hskip 28.45274pt\leq 3^{k}\tilde{d}(Z,U)^{-k}\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subseteq X\vee Y\end{subarray}}\tilde{d}(X,U)^{-4n-4}\tilde{d}(Y)^{-4n-4}$ (50)
It remains to bound the sum on the right-hand side. Fix an arbitrary $z\in Z$ and let $X,Y\in{\mathbb{B}}_{n}$ with $X\cap Y\neq\varnothing$ and $Z\subseteq X\vee Y$. Then by the second statement in Lemma 31 and the inequality $1+a+b\leq(1+a)(1+b)$ for $a,b\geq 0$ we have
@@
$\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subseteq X\vee Y\end{subarray}}\tilde{d}(X,U)^{-4n-4}\tilde{d}(Y)^{-4n-4}\\ \leq\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subseteq X\vee Y\end{subarray}}(1+\operatorname{diam}(X))^{-4n-4}(1+\operatorname{diam}(Y))^{-4n-4}\\ \leq\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\end{subarray}}(1+\operatorname{diam}(X)+d(X,z))^{-2n-2}(1+\operatorname{diam}(Y)+d(Y,z))^{-2n-2}\\ \leq\left(\sum_{\begin{subarray}{c}X\in{\mathbb{B}}_{n}\end{subarray}}(1+\operatorname{diam}(X)+d(X,z))^{-2n-2}\right)^{2}\leq\frac{\pi^{8}16^{n}(n+1)^{4}}{1296}.$ (52)
Now we are ready to prove Proposition 30.
Proof of Proposition 30.
Recall $\tilde{d}(X,U):=1+\operatorname{diam}(X)+d(X,U)$ and $\tilde{d}(X):=1+\operatorname{diam}(X)$.
Fix $Z\in{\mathbb{B}}_{n}$. By Lemma 32 we may restrict to pairs $(X,Y)$ with $X\cap Y\neq\varnothing$ and $Z\subseteq X\vee Y$. Moreover, by Lemma 11
$\|[{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]^{Z}\|\leq 4^{n}\|[{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]\|\leq 2^{2n+1}\|{\mathsf{F}}^{X}\|\,\|{\mathsf{G}}^{Y}\|.$ (53)
Using the definition of $\|\cdot\|_{U,\cdot}$ and $\|\cdot\|_{V,\cdot}$ we have
$\|{\mathsf{F}}^{X}\|$ $\leq\|{\mathsf{F}}\|_{U,k+4n+4}\,\tilde{d}(X,U)^{-k-4n-4}$ (54)
$\|{\mathsf{G}}^{Y}\|$ $\leq\|{\mathsf{G}}\|_{V,k+4n+4}\,\tilde{d}(Y,V)^{-k-4n-4}.$ (55)
Since $\tilde{d}(Y,V)=1+\operatorname{diam}(Y)+d(Y,V)\geq 1+\operatorname{diam}(Y)=\tilde{d}(Y)$, we also have $\tilde{d}(Y,V)^{-k-4n-4}\leq\tilde{d}(Y)^{-k-4n-4}$. Therefore
$\sum_{X,Y\in{\mathbb{B}}_{n}}\|[{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]^{Z}\|=\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subseteq X\vee Y\end{subarray}}\|[{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]^{Z}\|\\ \leq 2^{2n+1}\|{\mathsf{F}}\|_{U,k+4n+4}\|{\mathsf{G}}\|_{V,k+4n+4}\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ Z\subseteq X\vee Y\end{subarray}}\tilde{d}(X,U)^{-k-4n-4}\tilde{d}(Y)^{-k-4n-4}.$ (56)
By Lemma 33, the last sum is bounded by
$\frac{\pi^{8}16^{n}(n+1)^{4}}{1296}\,3^{k}\,\tilde{d}(Z,U)^{-k}.$ (57)
Hence
$\sum_{X,Y\in{\mathbb{B}}_{n}}\|[{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]^{Z}\|\leq\frac{\pi^{8}2^{6n+1}(n+1)^{4}}{648}\,3^{k}\,\|{\mathsf{F}}\|_{U,k+4n+4}\|{\mathsf{G}}\|_{V,k+4n+4}\,\tilde{d}(Z,U)^{-k}.$ (58)
In particular, the defining series (39) is absolutely convergent. Moreover,
$\|[{\mathsf{F}},{\mathsf{G}}]^{Z}\|\leq\sum_{X,Y}\|[{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]^{Z}\|\leq C\,3^{k}\,\|{\mathsf{F}}\|_{U,k+4n+4}\|{\mathsf{G}}\|_{V,k+4n+4}\,\tilde{d}(Z,U)^{-k}$ (59)
with $C=\frac{\pi^{8}2^{6n+1}(n+1)^{4}}{648}$. Multiplying by $\tilde{d}(Z,U)^{k}$ and taking the supremum over $Z$ yields (40).
Bilinearity and antisymmetry of the bracket are immediate. It remains to prove the Jacobi identity. It suffices to do it on ${\mathfrak{D}}_{al}({\mathbb{R}}^{n})$, since ${\mathfrak{D}}_{al}(U)\subseteq{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$. Let ${\mathsf{F}},{\mathsf{G}},{\mathsf{H}}\in{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ and fix $W\in{\mathbb{B}}_{n}$. Expanding ${\mathsf{G}},{\mathsf{H}}]^{Y}$ by definition gives the formal identity
${\mathsf{F}},[{\mathsf{G}},{\mathsf{H}}]]^{W}=\sum_{X,Y\in{\mathbb{B}}_{n}}\sum_{X^{\prime},Y^{\prime}\in{\mathbb{B}}_{n}}\Bigl[{\mathsf{F}}^{X},\,[{\mathsf{G}}^{X^{\prime}},{\mathsf{H}}^{Y^{\prime}}]^{Y}\Bigr]^{W}.$ (60)
We now show that the quadruple series in (60) is absolutely convergent. By Lemma 32, the term $\bigl[{\mathsf{F}}^{X},[{\mathsf{G}}^{X^{\prime}},{\mathsf{H}}^{Y^{\prime}}]^{Y}\bigr]^{W}$ can be nonzero only if
$X\cap Y\neq\varnothing,\quad W\subseteq X\vee Y,\qquad\text{and}\qquad X^{\prime}\cap Y^{\prime}\neq\varnothing,\quad Y\subseteq X^{\prime}\vee Y^{\prime}.$ (61)
Moreover, Lemma 11 implies
$\Bigl\|\bigl[{\mathsf{F}}^{X},[{\mathsf{G}}^{X^{\prime}},{\mathsf{H}}^{Y^{\prime}}]^{Y}\bigr]^{W}\Bigr\|\leq 4^{n}\Bigl\|\bigl[{\mathsf{F}}^{X},[{\mathsf{G}}^{X^{\prime}},{\mathsf{H}}^{Y^{\prime}}]^{Y}\bigr]\Bigr\|\leq 2^{2n+1}\|{\mathsf{F}}^{X}\|\,\|[{\mathsf{G}}^{X^{\prime}},{\mathsf{H}}^{Y^{\prime}}]^{Y}\|\\ \leq 2^{4n+2}\|{\mathsf{F}}^{X}\|\,\|{\mathsf{G}}^{X^{\prime}}\|\,\|{\mathsf{H}}^{Y^{\prime}}\|.$ (62)
Therefore
$\sum_{X,Y}\sum_{X^{\prime},Y^{\prime}}\Bigl\|\bigl[{\mathsf{F}}^{X},[{\mathsf{G}}^{X^{\prime}},{\mathsf{H}}^{Y^{\prime}}]^{Y}\bigr]^{W}\Bigr\|$
$\leq 2^{4n+2}\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ W\subseteq X\vee Y\end{subarray}}\|{\mathsf{F}}^{X}\|\sum_{\begin{subarray}{c}X^{\prime},Y^{\prime}\in{\mathbb{B}}_{n}\\ X^{\prime}\cap Y^{\prime}\neq\varnothing\\ Y\subseteq X^{\prime}\vee Y^{\prime}\end{subarray}}\|{\mathsf{G}}^{X^{\prime}}\|\,\|{\mathsf{H}}^{Y^{\prime}}\|.$ (63)
Bound the inner sum using $\|{\mathsf{G}}^{X^{\prime}}\|\leq\|{\mathsf{G}}\|_{{\mathbb{R}}^{n},8n+8}\tilde{d}(X^{\prime})^{-8n-8}$ and $\|{\mathsf{H}}^{Y^{\prime}}\|\leq\|{\mathsf{H}}\|_{{\mathbb{R}}^{n},8n+8}\tilde{d}(Y^{\prime})^{-8n-8}$, and apply Lemma 33 with $U={\mathbb{R}}^{n}$ and $k=4n+4$ to obtain
@@
$\sum_{\begin{subarray}{c}X^{\prime},Y^{\prime}\in{\mathbb{B}}_{n}\\ X^{\prime}\cap Y^{\prime}\neq\varnothing\\ Y\subseteq X^{\prime}\vee Y^{\prime}\end{subarray}}\|{\mathsf{G}}^{X^{\prime}}\|\,\|{\mathsf{H}}^{Y^{\prime}}\|\leq C_{1}\,\|{\mathsf{G}}\|_{{\mathbb{R}}^{n},8n+8}\|{\mathsf{H}}\|_{{\mathbb{R}}^{n},8n+8}\,\tilde{d}(Y)^{-4n-4},$ (64)
for a constant $C_{1}$ depending only on $n$. Plugging this into the outer sum and using $\|{\mathsf{F}}^{X}\|\leq\|{\mathsf{F}}\|_{{\mathbb{R}}^{n},4n+4}\tilde{d}(X)^{-4n-4}$ gives
@@
$\sum_{X,Y}\sum_{X^{\prime},Y^{\prime}}\Bigl\|\bigl[{\mathsf{F}}^{X},[{\mathsf{G}}^{X^{\prime}},{\mathsf{H}}^{Y^{\prime}}]^{Y}\bigr]^{W}\Bigr\|\leq\\ C_{2}\|{\mathsf{F}}\|_{{\mathbb{R}}^{n},4n+4}\|{\mathsf{G}}\|_{{\mathbb{R}}^{n},8n+8}\|{\mathsf{H}}\|_{{\mathbb{R}}^{n},8n+8}\sum_{\begin{subarray}{c}X,Y\in{\mathbb{B}}_{n}\\ X\cap Y\neq\varnothing\\ W\subseteq X\vee Y\end{subarray}}\tilde{d}(X)^{-4n-4}\tilde{d}(Y)^{-4n-4},$ (65)
and the remaining double sum is finite by Lemma 33 with $U={\mathbb{R}}^{n}$ and $k=0$. Hence the quadruple series (60) is absolutely convergent.
Absolute convergence justifies rearranging sums; in particular, since
$\sum_{Y\in{\mathbb{B}}_{n}}[{\mathsf{G}}^{X^{\prime}},{\mathsf{H}}^{Y^{\prime}}]^{Y}=[{\mathsf{G}}^{X^{\prime}},{\mathsf{H}}^{Y^{\prime}}],$ (66)
we obtain an absolutely convergent triple expansion
${\mathsf{F}},[{\mathsf{G}},{\mathsf{H}}]]^{W}=\sum_{X,Y,Z\in{\mathbb{B}}_{n}}[{\mathsf{F}}^{X},[{\mathsf{G}}^{Y},{\mathsf{H}}^{Z}]]^{W},$ (67)
and similarly for the cyclic permutations. For each fixed triple $(X,Y,Z)$, the ordinary commutator in the local algebra satisfies the Jacobi identity:
${\mathsf{F}}^{X},[{\mathsf{G}}^{Y},{\mathsf{H}}^{Z}]]+[{\mathsf{G}}^{Y},[{\mathsf{H}}^{Z},{\mathsf{F}}^{X}]]+[{\mathsf{H}}^{Z},[{\mathsf{F}}^{X},{\mathsf{G}}^{Y}]]=0.$ (68)
Taking the $W$-component and summing over $X,Y,Z$ gives the Jacobi identity for the bracket on ${\mathfrak{D}}_{al}({\mathbb{R}}^{n})$, hence on all ${\mathfrak{D}}_{al}(U)$. This completes the proof. ∎
From Propositions 28 and 30 we immediately get
Corollary 34.
Suppose $U,V\subseteq{\mathbb{R}}^{n}$.